OpenStudy (anonymous):

which equation represents the line passing through the points (3,2) and (-9,6) A. x-3y=9 B. x+3y=9 C. 3x-y=-9 D. 3x+y=9

Mehek (mehek14):

which one do you think?

OpenStudy (anonymous):

A

Mehek (mehek14):

it's not A because the line has (3,-2) not (3,2)

Mehek (mehek14):

but it has (-9,-6) not (-9,6)

OpenStudy (anonymous):

so i have to replace the point into the equation

Mehek (mehek14):

yes and see which one is true

Mehek (mehek14):

let's see B

Mehek (mehek14):

we will use the ordered pair (3,2) x = 3 y = 2

Mehek (mehek14):

so in the equation x+3y=9 substitute the values 3 + 3(2) = 9 3 + 3*2 = 9 is that true?

OpenStudy (anonymous):

yes

Mehek (mehek14):

ok now let's try the second ordered pair (-9,6) x+3y=9 -9 + 3(6) = 9 -9 + 3 * 6 = 9 true?

OpenStudy (anonymous):

yes

Mehek (mehek14):

so that is the equation for the line!

OpenStudy (anonymous):

thanks alot

Mehek (mehek14):

yw ^_^